OpenStudy (anonymous):

Use Newton's method to approximate a root of the equation 5x^3 +8x+1=0 as follows. Let x1 =−1 be the initial approximation. The second approximation x2 is ? and the third approximation x3 is ?

OpenStudy (anonymous):

The formula for newton's method is: \[x_{n+1} = x_{n} - f(x_{n})/f'(x_{n})\] if \[f(x) = 5x^3 + 8x + 1\] \[f'(x) = 15x^2 + 8\] So if \[x_{1} = -1\] then \[x_{2} = x_{1} - f(x_{1})/f'(x_{1})\] \[x_{2} = -1 - (5(-1)^3 + 8(-1) + 1)/(15(-1)^2 + 8)\] \[x_{2} \approx -0.4782\] then: \[x_{3} = x_{2} - f(x_{2})/f'(x_{2})\] When you work that out you'll get your answer for \[x_{3}\]

OpenStudy (anonymous):

excellent, thanks